As batteries discharge, their voltage decreases. The manufacturer of a cell phone battery would like to maintain 90.0% of its initial voltage. If a battery has an initial voltage of 5.95 V and loses an average of 0.040 V per minute, how long will it remain at 90.0% or more of the initial voltage? In minutes
Answer #1

Accumulation of Energy = inlet - outlet + concumption

no inlet/outlet so only consumption

Accumulation of Energy = consumption

V0 = 5.95 V

rate of loss = 0.04 V /min

time required to achieve 90% of 5.95 V?

find the 90% of that

0.9*5.95 = 5.355 V

0.1*5.95 = 0.595 V

so

time = V / rate

time = (0.595 V) /(0.04 V/min) = 14.875 minutes

in 14.875 minutes the charge will be at 90% of the initial+